OpenStudy (anonymous):

If x ÷ -1/6 = 0 , then x = ? 0 1 -1/6 1/6

OpenStudy (anonymous):

\[x \div 1/6=0\] \[x \times 6/1=0\] 6x=0 x=0/6=0